OpenStudy (anonymous):

Write the equation of a line in point-slope form that has a slope of and passes through the point (-6, 0).

OpenStudy (anonymous):

What is the slope given??

OpenStudy (anonymous):

y=mx+6

OpenStudy (anonymous):

Required equation is given by: \[\huge (y - y_1) = m(x - x_1)\]